OpenStudy (anonymous):

A doctor’s office schedules 15-minute appointments and half-hour appointments for weekdays. The doctor limits these appointments to, at most, 30 hours per week. Write an inequality to represent the number of 15-minute appointments x and the number of half-hour appointments y the doctor may have in a week.

OpenStudy (anonymous):

Hint: Time spent on 15 min appts + Time spent on 30 min appts <= 30 hours

OpenStudy (anonymous):

How can we write "time spent on 15 min appts" using a short algebraic expression?

OpenStudy (anonymous):

hello?................ if you want help on this website, you have to interact.

OpenStudy (anonymous):

Hey so I'm actually working on this same exact problem lol,.... Would it be 15x + 30y <= 30?

OpenStudy (anonymous):

Sorry I'm thinking t = 15 may be?

OpenStudy (anonymous):

Zezzy is on the right track. You'll need to convert that 30 (which is hours) to minutes since you have minutes on the left side of your inequality (15minutes and 30 minutes). Just multiply the 30 by 60

OpenStudy (anonymous):

1,800

OpenStudy (anonymous):

Ohhh right, 15x + 30y <= 180

OpenStudy (anonymous):

not 180... 1800

OpenStudy (anonymous):

Lol...Oops.... Brain isn't all the way awake I guess... X3 Thank you for the help.

OpenStudy (anonymous):

\(15x+30y\le 1800\) is the inequallity right? I feel dumb lol i used y = mx+b -.-

OpenStudy (anonymous):

Yeah I believe that's the answer they're looking for. :3